OpenStudy (anonymous):

Please help! (: Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 10 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

FIND Y IN TERM OF X FROM EQUATION 1

OpenStudy (anonymous):

huh...?

OpenStudy (anonymous):

PLUG IN THAT VALUE INTO SECOND EQUATION

OpenStudy (anonymous):

i dont know what you mean?

OpenStudy (anonymous):

Y EQUAL WHAT FROM FIRST EQUATION?

OpenStudy (anonymous):

y=10+x...?

OpenStudy (anonymous):

NO Y=X-10

OpenStudy (anonymous):

how..?

OpenStudy (anonymous):

ADD PLUS Y TO BOTH SIDE -10 FROM BOTH SIDE

OpenStudy (anonymous):

why are you typing in all caps??? and that makes no since...

OpenStudy (anonymous):



OpenStudy (anonymous):

wouldnt you solve it like this though? x-y=10 +y +y -------- x=10+y then substitute y into the second equation

OpenStudy (anonymous):

y=?

OpenStudy (anonymous):

y=-18?

OpenStudy (anonymous):

now y=x-10

OpenStudy (anonymous):

would x=28...?

OpenStudy (anonymous):

2x+x-10=2

OpenStudy (anonymous):

3x=12 x=4

OpenStudy (anonymous):

you confuse me :/

OpenStudy (anonymous):

get y from first equation plug in that y into the second thast it

OpenStudy (anonymous):

you still confuse me.... lol sorry

OpenStudy (anonymous):

x-y=10 what is y?

OpenStudy (anonymous):

i solved for x for the first equation....

OpenStudy (anonymous):

x-y=10 +y +y ------ x=10+y 2*10+y=2 20+y=2 -20 -20 -------- y=-18 2x-(-18)=10 +18 +18 ------------ 2x=28 /2 /2 ----- x=14 (14,-18)<==ANSWER right??